# Text Animation Illustration

### v2

@@ Abstract.
Text animation serves as an expressive medium, transforming static communication into dynamic experiences by infusing words with motion to evoke emotions, emphasize meanings, and construct compelling narratives. Crafting animations that are semantically aware poses significant challenges, demanding expertise in graphic design and animation. We present an automated text animation scheme, termed “Dynamic Typography”, which combines two challenging tasks. It deforms letters to convey semantic meaning and infuses them with vibrant movements based on user prompts. Our technique harnesses vector graphics representations and an end-to-end optimization-based framework. This framework employs neural displacement fields to convert letters into base shapes and applies per-frame motion, encouraging coherence with the intended textual concept. Shape preservation techniques and perceptual loss regularization are employed to maintain legibility and structural integrity throughout the animation process. We demonstrate the generalizability of our approach across various text-to-video models and highlight the superiority of our end-to-end methodology over baseline methods, which might comprise separate tasks. Through quantitative and qualitative evaluations, we demonstrate the effectiveness of our framework in generating coherent text animations that faithfully interpret user prompts while maintaining readability. Our code is available at: https://animate-your-word.github.io/demo/.
SVG, text-to-video generation, animation, kinetic typography
submissionid: 1234journal: FACMPjournal: TOG
1. Introduction
@@
Text animation is the art of bringing text to life through motion. By animating text to convey emotion, emphasize meaning, and create a dynamic narrative, text animation transforms static messages into vivid, interactive experiences (Lee et al., 2002b, 2006). The fusion of motion and text, not only captivates viewers, but also deepens the message’s impact, making text animation prevail in movies, advertisements, website widgets, and online memes (Xie et al., 2023).
@@
This paper introduces a specialized text animation scheme that focuses on animating individual letters within words. This animation is a compound task: The letters are deformed to embody their semantic meaning and then brought to life with vivid movements based on the user’s prompt. We refer to it as “Dynamic Typography”. For example, the letter “M” in “CAMEL” can be animated with the prompt “A camel walks steadily across the desert” as illustrated in Fig. 1. This animation scheme opens up a new dimension of textual animation that enriches the user’s reading experience.
@@
However, crafting such detailed and prompt-aware animations is challenging, as traditional text animation methods demand considerable expertise in graphic design and animation (Lee et al., 2002b), making them less accessible to non-experts. The technique we present aims to automate the text animation process to make it more accessible and efficient. Following prior research in font generation and stylization (Lopes et al., 2019; Wang and Lian, 2021; Iluz et al., 2023), we represent each input letter and every output frame as a vectorized, closed shape by a collection of Bézier curves. This vector representation is resolution-independent, ensuring that text remains clear and sharp regardless of scale, and brings substantial editability benefits as users can easily modify the text’s appearance by adjusting control points. However, this shift to vector graphics introduces unique challenges in text animation. Most current image-to-video methods (Ni et al., 2023; Wang et al., 2024; Xing et al., 2023; Guo et al., 2024) fall short in this new scenario as they are designed to animate rasterized images instead of vectorized shapes, and are hard to render readable text. Although the most recent work, LiveSketch (Gal et al., 2023), introduces an approach to animate arbitrary vectorized sketches, it struggles to preserve legibility and consistency during animation when the input becomes vectorized letters, causing visually unpleasant artifacts including flickering and distortion.
@@
To address these challenges, we designed an optimization-based end-to-end framework that utilizes two neural displacement fields, represented in coordinates-based MLP. The first field deforms the original letter into the base shape, setting the stage for animation. Subsequently, the second neural displacement field learns the per-frame motion applied to the base shape. The two fields are jointly optimized using the score-distillation sampling (SDS) loss (Poole et al., 2023) to integrate motion priors from a pre-trained text-to-video model (Wang et al., 2023), to encourage the animation to align with the intended textual concept. We encode the control point coordinates of the Bézier curve into high-frequency encoding (Mildenhall et al., 2021) and adopt coarse-to-fine frequency annealing (Park et al., 2021) to capture both minute and large motions. To preserve the legibility of the letter throughout the animation, we apply perceptual loss (Zhang et al., 2018) as a form of regularization on the base shape, to maintain a perceptual resemblance to the original letter. Additionally, to preserve the overall structure and appearance during animation, we introduce a novel shape preservation regularization based on the triangulation (Hormann and Greiner, 2000) of the base shape, which forces the deformation between the consecutive frames to adhere to the principle of being conformal with respect to the base shape.
@@
Our approach is designed to be data-efficient, eliminating the need for additional data collection or the fine-tuning of large-scale models. Furthermore, our method generalizes well to various text-to-video models, enabling the incorporation of upcoming developments in this area. We quantitatively and qualitatively tested our text animation generation method against various baseline approaches, using a broad spectrum of prompts. The results demonstrate that the generated animation not only accurately and aesthetically interprets the input text prompt descriptions, but also maintains the readability of the original text, outperforming various baseline models in preserving legibility and prompt-video alignment. Overall, our framework demonstrates its efficacy in producing coherent text animations from user prompts, while maintaining the readability of the text, which is achieved by the key design of the learnable base shape and associated shape preservation regularization.
2. Related Work
2.1. Static Text Stylization
Text stylization focuses on amplifying the aesthetic qualities of text while maintaining readability, including artistic text style transfer and semantic typography. Artistic text style transfer aims to migrate stylistic elements from a source image onto text. Existing work incorporates texture synthesis (Fish et al., 2020; Yang et al., 2016) with generative models like GANs (Azadi et al., 2018; Jiang et al., 2019; Mao et al., 2022; Wang et al., 2019). Semantic typography refers to techniques that blend semantic understanding and visual representation in typography. This encompasses turning letters or words into visual forms that convey their meaning or nature, integrating typography with semantics to enhance the message’s clarity and impact. For instance, (Iluz et al., 2023) leverages Score Distillation Sampling (Poole et al., 2023) to deform letters based on the pre-trained diffusion prior (Rombach et al., 2022), encouraging the appearance of the letter to convey the word’s semantic meaning. (Tanveer et al., 2023) utilizes a latent diffusion process to construct the latent space of the given semantic-related style and then introduces a discriminator to blend the style into the glyph shape.
These works only produce static images, which in many cases struggle to vividly and effectively communicate meaningful semantic messages. In contrast, our proposed “Dynamic Typograph” infuses text with vibrant motions, which is more effective in capturing the user’s attention and giving an aesthetically pleasing impression compared to static text (Minakuchi and Kidawara, 2008).
2.2. Dynamic Text Animation
Given the effectiveness of animations in capturing and retaining audience attention (Chang and Ungar, 1993), several studies have embarked on designing dynamic text animations. A notable area is dynamic style transfer, which aims to adapt the visual style and motion patterns from a reference video to the target text. Pioneering work by (Men et al., 2019) transferred a style from a source video displaying dynamic text animations onto target static text. (Yang et al., 2021) further enhanced versatility by using a scale-aware Shape-Matching GAN to handle diverse input styles.
Kinetic typography (Ford et al., 1997) represents another innovative direction in text animation, which integrates motion with text to convey or enhance a message. Creating kinetic typography is a labor-intensive and challenging task, motivating many works to reduce the burden and democratize this technique to the public (Lee et al., 2002a; Ford et al., 1997; Minakuchi and Tanaka, 2005; Forlizzi et al., 2003). Recent advancements in AI have enabled a fully automated generation of kinetic typography. For example, (Xie et al., 2023) utilizes a pre-trained motion transfer model (Siarohin et al., 2019) to apply animation patterns from a meme GIF onto text.
However, these approaches require strictly specified driven videos, which are difficult to obtain in real-life scenarios, significantly restricting their usability and generalizability. Moreover, they are constrained to generate specific simple motion patterns, limiting their ability to produce animations with arbitrary complex semantic information. In contrast, our method is generalizable to arbitrary motion patterns and only needs a text prompt as the input.
@@ 2.2. Dynamic Text Animation
Figure 2. An overview of the model architecture. Given a letter represented as a set of control points, the Base Field deforms it to the shared base shape, setting the stage to add per-frame displacement. Then we duplicate the base shape across $k$ frames and utilize the Motion Field to predict displacements for each control point at each frame, infusing movement to the base shape. Every frame is then rendered by the differentiable rasterizer $R$ and concatenated as the output video. The base and motion field are jointly optimized by the video prior ($L_{SDS}$) from frozen pre-trained video foundation model using Score Distillation Sampling, under regularization on legibility $L_{legibility}$ and structure preservation $L_{structure}$.
2.3. Text and Image-to-Video Generation
Text-to-Video generation aims at automatically producing corresponding videos based on textual descriptions. Recent advancements in diffusion models have significantly improved video generation capabilities. Mainstream approaches leverage the power of Stable Diffusion (SD) (Rombach et al., 2022) by incorporating temporal information in a latent space, including AnimateDiff (Guo et al., 2024), LVDM (He et al., 2022), MagicVideo (Zhou et al., 2022), VideoCrafter (Chen et al., 2023) and ModelScope (Wang et al., 2023). Beyond text-to-video, some methods attempt to generate videos from a given image and a prompt as the condition, such as DynamiCrafter (Xing et al., 2023) and Motion-I2V (Shi et al., 2024). Several startups also release their image-to-video generation services, e.g., Gen-2 (contributors, 2023a), Pika Labs (contributors, 2023b), and Stable Video Diffusion (SVD) (Blattmann et al., 2023).
@@ 2.3. Text and Image-to-Video Generation
Despite progress, open-source video generation models struggle to maintain text readability during motion, let alone create vivid text animations. Training a model capable of generating high-quality, legible text animations using the aforementioned methods would require a large dataset of text animations, which is difficult to require in practice. One recent work LiveSketch (Gal et al., 2023) introduces an approach to animate arbitrary vectorized sketches without extensive training. This work leverages the motion prior from a large pre-trained text-to-video diffusion model using score distillation sampling (Poole et al., 2023) to guide the motion of input sketches. However, when the input becomes vectorized letters, LiveSketch struggles to preserve legibility and consistency during animation, leading to flickering and distortion artifacts that severely degrade video quality. In contrast, our proposed method successfully generates consistent and prompt-aware text animations while preserving the text readability.
3. Preliminary
3.1. Vector Representation and Fonts
Vector graphics create visual images directly from geometric shapes like points, lines, curves, and polygons. Unlike raster images (like PNG and JPEG), which store data for each pixel, vector graphics are not tied to a specific resolution, making them infinitely scalable and more editable (Ferraiolo et al., 2000).
@@ 3.1. Vector Representation and Fonts
Hence, modern font formats like TrueType (Penny, 1996) and PostScript (Adobe Systems Inc., 1990) utilize vector graphics to define glyph outlines. These outlines are typically collections of Bézier or B-Spline curves, enabling scalable and flexible text rendering, which we aim to preserve. Our method outputs each animation frame in the same vector representations as our input.
Figure 3. Bźeier curves representation of letter “B”
@@
In alignment with the setting outlined in (Iluz et al., 2023), we use the FreeType (David Turner, 2009) font library to extract the outlines of the specified letter. Subsequently, these outlines are converted into a closed curve composed of several cubic Bézier curves, as illustrated in Fig. 3, to achieve a coherent representation across different fonts and letters.
3.2. Score Distillation Sampling
@@
The objective of Score Distillation Sampling (SDS), originally introduced in the DreamFusion (Poole et al., 2023), is to leverage pre-trained diffusion models’ prior knowledge for the text-conditioned generation of different modalities (Katzir et al., 2024). SDS optimizes the parameters $\theta$ of the parametric generator $\mathcal{G}$ (e.g., NeRF (Mildenhall et al., 2021)), ensuring the output of $\mathcal{G}$ aligns well with the prompt. For illustration, assuming $\mathcal{G}$ is a parametric image generator. First, an image $x=\mathcal{G}(\theta)$ is generated. Next, a noise image $z_{\tau}(x)$ is obtained by adding a Gaussian noise $\epsilon$ at the diffusion process’s $\tau$-th timestep:
(1) $z_{\tau}(x)=\alpha_{\tau}x+\sigma_{\tau}\epsilon,$
@@
where $\alpha_{\tau}$, and $\sigma_{\tau}$ are diffusion model’s noising schedule, and $\epsilon$ is a noise sample from the normal distribution $\mathcal{N}(0,1)$.
@@
(2) $\nabla_{\phi}\mathcal{L}_{SDS}=\left[w(\tau)(\epsilon_{\phi}(z_{\tau}(x);y,\tau)-\epsilon)\frac{\partial x}{\partial\theta}\right],$
@@
where $y$ is the conditioning input to the diffusion model and $w(\tau)$ is a weighting function. The diffusion model predicts the noise added to the image $x$ with $\epsilon_{\phi}(z_{\tau}(x);y,\tau)$. The discrepancy between this prediction and the actual noise $\epsilon$ measures the difference between the input image and one that aligns with the text prompt. In this work, we adopt this strategy to extract the motion prior from the pre-trained text-to-video diffusion model (Wang et al., 2023).
@@
Since SDS is used with raster images, we utilize DiffVG (Li et al., 2020) as a differentiable rasterizer. This allows us to convert our vector-defined content into pixel space in a differentiable way for applying the SDS loss.
4. Method
@@
Problem Formulation. Dynamic Typography focuses on animating individual letters within words based on the user’s prompt. The letter is deformed to embody the word’s semantic meaning and then brought to life by infusing motion based on the user’s prompt.
@@
The original input letter is initialized as a cubic Bézier curves control points set (Fig. 3), denoted as $P_{letter}=\{p_{i}\}_{i=1}^{N}=\{(x_{i},y_{i})\}_{i=1}^{N}\in\mathbb{R}^{N\times 2}$, where $x,y$ refers to control points’ coordinates in SVG canvas, $N$ refers to the total number of control points of the indicated letter. The output video consists of k frames, each represented by a set of control points, denoted as $V=\{P^{t}\}_{t=1}^{k}\in\mathbb{R}^{N\cdot k\times 2}$, where $P^{t}$ is the control points for $t$-th frame.
@@
Our goal is to learn a displacement for each frame, added on the set of control point coordinates of the original letter’s outline. This displacement represents the motion of the control points over time, creating the animation that depicts the user’s prompt. We denote the displacement for $t$-th frame as $\Delta P^{t}=\{\Delta p_{i}^{t}\}_{i=1}^{N}=\{(\Delta x_{i}^{t},\Delta y_{i}^{t})\}_{i=1}^{N}\in\mathbb{R}^{N\times 2}$,where $\Delta p_{i}^{t}$ refers to the displacement of the $i$-th control point in the $t$-th frame. The final video can be derived as $V=\{P_{letter}+\Delta P^{t}\}_{t=1}^{k}$.
To achieve appealing results, we identify three crucial requirements for Dynamic Typography: (1) Temporal Consistency. The deformed letter should move coherently while preserving a relatively consistent appearance in each animation frame. (2) Legibility Preservation. The deformed letter should remain legible in each frame during animation. (3) Semantic Alignment. The letter should be deformed and animated in a way that aligns with the semantic information in the text prompt.
@@
One straightforward strategy can be first deforming the static letter with existing methods like (Iluz et al., 2023), then utilizing an animation model designed for arbitrary graphics composed of Beźier curves like (Gal et al., 2023) to animate the deformed letter. However, this non-end-to-end formulation suffers from conflicting prior knowledge. The deformed letter generated by the first model may not align with the prior knowledge of the animation model. This mismatch can lead the animation model to alter the appearance of the deformed letter, leading to considerable visual artifacts including distortion and inconsistency, see Fig. 4.
@@
Figure 4. Illustration of the prior knowledge conflict issue. The left is the deformed “R” for BULLFIGHTER with prompt “A bullfighter holds the corners of a red cape in both hands and waves it” generated by (Iluz et al., 2023), the right is generated by (Gal et al., 2023) to animate the deformed letter with the same prompt. The mismatch in prior knowledge between separate models leads to significant appearance changes and severe artifacts, as highlighted by the red circles.
@@
Therefore, to ensure the coherence of the entire process, we propose an end-to-end architecture that directly maps the original letter to the final animation, as illustrated in Fig. 2. To address the complexity of learning per-frame displacement that converts the input letter into animation, we represent the video as a learnable base shape and per-frame motion added on the base shape (§4.1). Additionally, we incorporate legibility regularization based on perceptual similarity to maintain letter legibility (§4.2). Then, we introduce a mesh-based structure preservation loss to ensure appearance and structure integrity between frames, mitigating issues such as flickering artifacts (§4.3). Finally, we utilize frequency-based encoding and coarse-to-fine annealing to improve the representation of geometry information and motion quality (§4.4).
4.1. Base Field and Motion Field

### v3

# Dynamic Typography: Bringing Text to Life via Video Diffusion Prior

@@ Abstract
Text animation serves as an expressive medium, transforming static communication into dynamic experiences by infusing words with motion to evoke emotions, emphasize meanings, and construct compelling narratives. Crafting animations that are semantically aware poses significant challenges, demanding expertise in graphic design and animation. We present an automated text animation scheme, termed “Dynamic Typography”, which combines two challenging tasks. It deforms letters to convey semantic meaning and infuses them with vibrant movements based on user prompts. Our technique harnesses vector graphics and an end-to-end optimization-based framework. This framework employs neural displacement fields to convert letters into base shapes and applies per-frame motion, encouraging coherence with the intended textual concept. Perceptual loss regularization and shape preservation techniques are employed to maintain legibility and structural integrity throughout the animation process. We demonstrate the generalizability of our approach across various text-to-video models and highlight the superiority of our end-to-end methodology over baseline methods, which might comprise separate tasks. Through quantitative and qualitative evaluations, we demonstrate the effectiveness of our framework in generating coherent text animations that faithfully interpret user prompts while maintaining readability.
Figure 1: Given a letter and a text prompt that briefly describes the animation, our method automatically semantically reshapes a letter and animates it in vector format while maintaining legibility. Our approach allows for a variety of creative interpretations that can dynamically bring words to life. Our code is available at: https://animate-your-word.github.io/demo/.
**footnotetext: Indicates Equal Contribution.$\dagger$\$$\dagger$\$footnotetext: Indicates Corresponding Author.
1 Introduction
@@
Text animation is the art of bringing text to life through motion. By animating text to convey emotion, emphasize meaning, and create a dynamic narrative, text animation transforms static messages into vivid, interactive experiences [25, 23]. The fusion of motion and text, not only captivates viewers, but also deepens the message’s impact, making text animation prevail in movies, advertisements, website widgets, and online memes [56].
@@
This paper introduces a specialized text animation scheme that focuses on animating individual letters within words. This animation is a compound task: The letters are deformed to embody their semantic meaning and then brought to life with vivid movements based on the user’s prompt. We refer to it as “Dynamic Typography”. For example, the letter “M” in “CAMEL” can be animated with the prompt “A camel walks steadily across the desert” in Fig. 1. This animation scheme opens up a new dimension of textual animation that enriches the reading experience.
@@
However, crafting such detailed and prompt-aware animations is challenging. Traditional text animation methods demand considerable expertise in graphic design and animation [25], making them less accessible to non-experts. Our methodology aims to automate the text animation process to make it more accessible and efficient. Following prior research in font generation and stylization [28, 54, 19], we represent each input letter and every output frame as a vectorized, closed shape by a collection of Bézier curves. This vector representation is resolution-independent, keeping text clear at any scale and offering substantial editability, as users can easily adjust text appearance through control points. However, this shift to vector graphics introduces unique challenges in text animation. Most current video generation methods [36, 53, 7, 57, 15] fall short in this new scenario as they are designed within the rasterized pixel-based scenario instead of vectorized shapes, and are hard to render readable text. Although the most recent work, LiveSketch [14], introduces an approach to animate arbitrary vectorized sketches, it struggles to preserve legibility and consistency throughout animation when the input becomes vectorized letters, causing visually unpleasant artifacts including flickering and distortion.
@@
To address these challenges, we design an end-to-end optimization-based framework that utilizes two neural displacement fields, represented in coordinates-based MLP. The first field deforms the original letter into the base shape, setting the stage for animation. Subsequently, the second neural displacement field learns the per-frame motion applied to the base shape. The two fields are jointly optimized using the score-distillation sampling (SDS) loss [42] to integrate motion priors from a pre-trained text-to-video model [51] to encourage the animation to align with the intended textual concept. To preserve the legibility of the letter throughout the animation, we apply perceptual loss [60] as a form of regularization on the base shape to maintain a perceptual resemblance to the original letter. Additionally, to preserve the overall structure and appearance during animation, we introduce a novel shape preservation regularization based on the triangulation [17] of the base shape, which forces the deformation between the consecutive frames to adhere to the principle of being conformal with respect to the base shape.
@@
Our approach is designed to be data-efficient, eliminating the need for additional data collection or the fine-tuning of large-scale models. Furthermore, our method generalizes well to various text-to-video models, enabling the incorporation of upcoming developments in this area. We quantitatively and qualitatively tested our text animation generation method against various baseline approaches, using a broad spectrum of prompts. The results demonstrate that the generated animation accurately and aesthetically interprets the input text prompt descriptions. Our method outperforms various baseline models in preserving legibility and prompt-video alignment. Overall, our framework demonstrates its efficacy in producing coherent text animations from user prompts while maintaining the readability of the text, which is achieved by the key design of the learnable base shape and associated shape preservation regularization.
2 Related Work
2.1 Static Text Stylization
Text stylization enhances text aesthetics while maintaining readability, including artistic text style transfer and semantic typography. Artistic text style transfer migrates stylistic elements from source images onto text, typically using texture synthesis [10, 58] and GANs [3, 20, 31, 52]. Semantic typography combines semantic understanding with visual representation, creating visual forms that convey meaning. Notable works include Word-as-Image [19], which uses Score Distillation Sampling [42] with diffusion prior [45] to create meaningful letter deformations, and DS-Fusion [49], which employs latent diffusion to blend semantic-related styles into glyphs.
While these works produce only static images with limited semantic expression, our Dynamic Typography introduces motion to text, enhancing viewer engagement and aesthetic appeal [34].
2.2 Dynamic Text Animation
Given animations’ effectiveness in capturing attention [6], research has explored dynamic text animations, particularly in dynamic style transfer and kinetic typography. Dynamic style transfer adapts visual style and motion from reference videos to text, with works like [32] transferring animations from dynamic text videos, and Yang et al. [59] introducing a scale-aware Shape-Matching GAN for diverse styles.
Kinetic typography [12] integrates motion with text to enhance messages. While traditionally labor-intensive, recent works [24, 12, 35, 13, 55] aim to automate this process. For instance, Wakey-Wakey [56] uses a motion transfer model [48] to animate text using meme GIFs.
However, these approaches require specific driven videos and are limited to simple motion patterns. In contrast, our method generates arbitrary motion patterns using only text prompts as input.
@@ 2.2 Dynamic Text Animation
Figure 2: An overview of the framework. Given a letter represented as a set of control points, the Base Field deforms it to the shared base shape, setting the stage to add per-frame displacement. Then the base shape is duplicated across $k$ frames, and the Motion Field predicts the displacement for each control point at each frame, infusing movement into the base shape. Each frame is rendered by the differentiable rasterizer $R$ and concatenated as the output video. The base and motion field are jointly optimized by the video prior from frozen pre-trained video foundation model using Score Distillation Sampling $\mathcal{L}_{\text{SDS}}$, under regularization on legibility $\mathcal{L}_{\text{legibility}}$ and structure preservation $\mathcal{L}_{\text{structure}}$.
2.3 Text and Image-to-Video Generation
Text-to-Video generation has advanced significantly with diffusion models. Several approaches extend Stable Diffusion (SD) [45] by incorporating temporal information, including AnimateDiff [15], LVDM [16], MagicVideo [61], VideoCrafter [7], and ModelScope [51]. Image-to-Video generation methods like DynamiCrafter [57], Motion-I2V [47], Gen-2 [46], Pika Labs [22], and SVD [5] generate videos from images and prompts.
@@ 2.3 Text and Image-to-Video Generation
However, existing open-source models fail to maintain text readability during motion. Training models for text animation requires large text animation datasets, which are scarce. One recent work LiveSketch [14] animates arbitrary vectorized sketches without extensive training. This work leverages the motion prior from pre-trained text-to-video diffusion model using score distillation sampling [42] to guide the motion of input sketches. However, when the input becomes vectorized letters, LiveSketch struggles to preserve legibility and consistency throughout the animation, leading to flickering and distortion artifacts that severely degrade video quality. Our method, in contrast, generates consistent, prompt-aware text animations while preserving readability.
3 Preliminary
3.1 Vector Representation and Fonts
@@
Modern font formats like TrueType [41] and PostScript [2] utilize vector graphics to define glyph outlines. These outlines are typically collections of Bézier or B-Spline curves, enabling flexible text rendering at any scale [19]. Our method outputs each animation frame in the same vector representation as our input.
Figure 3: Bézier curves representation of letter “B”. The endpoints are marked in orange, and the inner control points are in blue.
@@
In alignment with the setting outlined in Iluz et al. [19], we use the FreeType [8] font library to extract the outlines of the specified letter. Subsequently, these outlines are converted into a closed curve composed of several cubic Bézier curves, as illustrated in Fig. 3, to achieve a coherent representation across different fonts and letters. We iteratively subdivide the letter’s Bézier segments until reaching a pre-defined threshold to ensure sufficient control points to represent semantic deformation.
3.2 Score Distillation Sampling
@@
The objective of Score Distillation Sampling (SDS), originally introduced in the DreamFusion [42], is to leverage pre-trained diffusion models’ prior knowledge for the text-conditioned generation of different modalities [21]. SDS optimizes the parameters $\theta$ of the parametric generator $\mathcal{G}$ (e.g., NeRF [33]), ensuring the output of $\mathcal{G}$ aligns well with the prompt. For illustration, assuming $\mathcal{G}$ is a parametric image generator. First, an image $x=\mathcal{G}(\theta)$ is generated. Next, a noise image $z_{\tau}(x)$ is obtained by adding a Gaussian noise $\epsilon$ at the diffusion process’s $\tau$-th timestep:
$z_{\tau}(x)=\alpha_{\tau}x+\sigma_{\tau}\epsilon,$ (1)
@@
where $\alpha_{\tau}$, and $\sigma_{\tau}$ are diffusion model’s noising schedule, and $\epsilon$ is a noise sample from the distribution $\mathcal{N}(0,1)$.
@@
$\nabla_{\phi}\mathcal{L}_{SDS}=\left[w(\tau)(\epsilon_{\phi}(z_{\tau}(x);y,\tau)-\epsilon)\frac{\partial x}{\partial\theta}\right],$ (2)
@@
where $y$ is the conditioning input to the diffusion model and $w(\tau)$ is a weighting function. The diffusion model predicts the noise added to the image $x$ with $\epsilon_{\phi}(z_{\tau}(x);y,\tau)$. The discrepancy between this prediction and the actual noise $\epsilon$ measures the difference between the input image and one that aligns with the text prompt. In this work, we adopt this strategy to extract the motion prior from the pre-trained text-to-video diffusion model [51].
@@
Since SDS is used within the pixel-based scenario, we utilize DiffVG [26] as a differentiable rasterizer. This allows us to convert our vector-defined content into pixel space in a differentiable way for applying the SDS loss.
4 Method
@@
Dynamic Typography focuses on animating individual letters within words based on the user’s prompt. The letter is deformed to embody the word’s semantic meaning and then brought to life by infusing motion based on the user’s prompt. To achieve visually appealing animations, we identify three crucial requirements for Dynamic Typography:
  • Legibility Preservation. The deformed letter should remain legible in each frame during animation.
  • Semantic Alignment. The letter should be deformed and animated in a way that aligns with the semantic information in the text prompt.
  • Temporal Consistency. The deformed letter should move coherently while preserving a relatively consistent appearance in each animation frame.
@@
Problem Formulation: The original input letter is initialized as a cubic Bézier curves control points set (Fig. 3), denoted as $P_{letter}=\{p_{i}\}_{i=1}^{N}=\{(x_{i},y_{i})\}_{i=1}^{N}\in\mathbb{R}^{N\times 2}$, where $x,y$ refers to control points’ coordinates in SVG canvas, and $N$ is the total number of control points of the indicated letter. The output video consists of k frames, each represented by a set of control points, denoted as $V=\{P^{t}\}_{t=1}^{k}\in\mathbb{R}^{k\times N\times 2}$, where $P^{t}$ is the control points for $t$-th frame.
@@
Our goal is to learn the per-frame displacement to be added on the set of control point coordinates of the original letter’s outline. This displacement represents the motion of the control points over time, creating the animation that depicts the user’s prompt. We denote the displacement for $t$-th frame as $\Delta P^{t}=\{\Delta p_{i}^{t}\}_{i=1}^{N}=\{(\Delta x_{i}^{t},\Delta y_{i}^{t})\}_{i=1}^{N}\in\mathbb{R}^{N\times 2}$, where $\Delta p_{i}^{t}$ refers to the displacement of the $i$-th control point in the $t$-th frame. The final video can be derived as $V=\{P_{letter}+\Delta P^{t}\}_{t=1}^{k}$.
@@
One straightforward strategy can be first deforming the static letter with existing method [19], then utilizing an animation model [14] designed for vector graphics composed of Bézier curves to animate the deformed letter. However, this non-end-to-end formulation suffers from conflicting prior knowledge. The deformed letter generated by the first model may not align with the prior knowledge of the animation model. This mismatch can lead the animation model to alter the appearance of the deformed letter, leading to considerable visual artifacts including distortion and inconsistency, see Fig. 4.
@@
Figure 4: Illustration of the prior knowledge conflict issue. The left is the deformed “R” for “BULLFIGHTER” with prompt “A bullfighter holds the corners of a red cape in both hands and waves it” generated by Iluz et al. [19], the right is generated by Gal et al. [14] to animate the deformed letter with the same prompt. The mismatch in prior knowledge between separate models leads to significant appearance changes and severe artifacts, as highlighted by the red circles.
@@
Therefore, to ensure the coherence of the entire process, we propose an end-to-end framework that directly maps the original letter to the final animation, as illustrated in Fig. 2. To address the complexity of learning per-frame displacement that converts the letter into animation, we represent the video as a learnable base shape and per-frame motion added on the base shape (§4.1). Additionally, we incorporate legibility regularization loss based on perceptual similarity to maintain letter legibility (§4.2). Finally, we introduce a mesh-based structure preservation regularization loss to ensure appearance and structure integrity between frames, mitigating inconsistent artifacts (§4.3).
4.1 Base Field and Motion Field